Skip to end of metadata
Go to start of metadata

For the quickest system and user support of hardware and software, email the computing group at Requests are immediately forwarded to appropriate staff. See Helpdesk 101 for details, and to speed resolution of requests.  Clicking on "Core Services Status" below, will take you to a page with more detail on current system outages.


Core Services Status

CALENDAR - Department Calendar Services.

CERBERUS - /opt/local and shared data services.

DHCP - Wired DHCP services for the department.

DOCUSHARE - DocuShare database.

FTP - Anonymous FTP services.

HYDRA - HET Observing Support

IDL - IDL License server

MAIL - Astro mail services

MATHEMATICA - Mathematica License server

PRINTERS - Department Printer services

WEBMAIL - Webmail services

WEBSEVERS - Department websites and related services

HELPDESK - Helpdesk server


Individual Printer Status

If Printers under Core Services Status is Green, then the printer queue is active.  Below lists the public printers and their individual status in case one of them fail.

COPIERS - All Ricoh copiers

DW - Department color

LASER13 - 13th floor black and white

LASER15 - 15th floor black and white

LASER16 - 16th floor black and white



Current Public FTP Outage

  • 29 Mar @ 12:30 PM Anonymous FTP services restored.  
  • 24 Mar @ 8:31 PM Anonymous FTP services are currently unavailable due to file system problem with the /ftp/pub mount.
    • Unscheduled maintenance is planned for Monday, Mar 27. 

Scheduled Maintenance

  • 22 NO  @12:08 All systems updated and back online
  • 22 NO  @10:30 All systems updated and back online EXCEPT affirmed, which hosts the web sites and Mathematica License server
  • 22 NOV @7:29 - All major systems off line
  • 22 NOV @7:00-12:00- All majors systems will be off line

Scheduled Maintenance

  • 26 Sept @11:45 AM - DocuShare server will be taken off-line to install a new License file and clone the hard drive
  • 26 Sept @  3:31 PM - Disk cloning is going very,  very slowly. We hope to have Docushare back on-line until sometime tomorrow.

Scheduled Maintenance

  • 5 AUG @12:00 PM - Server reboot for astro, gemini, ftp, bigbubba to applied Oracle quarterly patches for Solaris OS - The following services will be affected:
    • email - incoming, outgoing, pine, and
    • Anonymous FTP
    • NFS exports for /san/pgs, /san/jk1, /san/jk2, /san/ne2, /san/ne3, /san/dw

    • logins to astro and bigbubba 

  • Still having trouble between astro and our main RAID....

  • 5 AUG @5:15 PM All services restoresd.

Scheduled Maintenance

  • 12 July @12:00 PM - Server reboot for astro, gemini, ftp, bigbubba to applied Oracle quarterly patches for Solaris OS - The following services will be affected:
    • email - incoming, outgoing, pine, and
    • Anonymous FTP
    • NFS exports for /san/pgs, /san/jk1, /san/jk2, /san/ne2, /san/ne3, /san/dw

    • logins to astro and bigbubba

  • All services  restored by 12:30 PM. 

Scheduled Maintenance

  • 31 March   @12:00 PM - Mail will be restarted.
  • 31 March   @12:23 PM - All access to email has been restored.

Scheduled Maintenance

  • 15 March   @11:45 AM - All access to email via TLSv1 has been restored.
  • 14 March   @ 11:00 AM - We have detected an issue with the Apple Mail applications IMAP connecting in with the existing settings after the SSL change.  For those users, please VPN and use Astro Webmail: until the fix is completed.
  • 14 March   @ 6:45 AM - Mail will be restarted

Scheduled Maintenance

  • 9 February @  12:09 - All work completed.
  • 9 February @ 7:17 - rsync started
  • 9 February @ 6:30 - move astro home directories from the 6140 RAID to cerberus, our Oracle 7120 RAID due to failed interface controller. All astro, web, mail, webmail services will be unavailable during this time. 

Unscheduled Maintenance

  • 21 December @10:20 AM - email services restored. 
  • 21 December @10:00 AM - astro email services will be stopped and restarted to clear hung/orphaned processes. Webmail will not be available during this period. Services are expected to back online by 10:20AM.


Scheduled Maintenance

  • 12 December @ 01:05 PM web sites passed basic testing (i.e., mounts from astro including home web pages work fine)
  • 12 December @ 12:24 PM Anonymous FTP test failed. did not remount the /ftp/pub share from astro on reboot. Manually mounted and successfully tested.
  • 12 December @ 12:23 PM test emails start showing up. 
  • 12 December @ 11:57 AM email testing began. 
  • 12 December @ 11:56 AM astro, ftp, bigbubba, solmaster, and gzone03, 
  • 12 December @ 11:56 AM gemini rebooted successfully with recommended and security patched applied
  • 12 December @ 10:39 AM - Zones BEs conflicted with global updates, requiring pkg update to be re-run on gemini (global zone).
  • 12 December @ 9:41 AM - Having trouble with some of the zones' boot environments (BE), but moving forward. Sorry for the delay 
  • 12 December @ 9:00 AM Most of our Solaris systems will be rebooted to apply updates that include security patches and new firewall features that we need to avoid quarantines by ISO. The following systems and services will be unavailable and offline for hopefully no longer than 20 minutes.
    • astro  - logins, all email services including webmail, webpages based located in astro home directories
    • gemini - exported RAID files systems /san/dw (Mike Montgomery (, /san/ne2  and /san/ne3 (Neal Evans' hick machines) /san/psg ( 
    • ftp - anonymous ftp
    • galactica - should have no user impact
    • gandhi - may take longer than 20 minutes to deal with Solaris 10 OS updates, which take longer than Solaris 11

  • 2 December @ 7:43 am all system are back online. The update to Cerberus was successful. Please report any anomalies to
  • 2 December @ 7:00 AM - One of our departmental file servers, Cerberus, is going to require a reboot to update the systems software. This update is required so we can install a patch to the system. The patch may require a reboot as well. We will know as soon as the Oracle engineers evaluate a resolution to a potential security issues identified by ISO. Use this wiki page to check status of this maintenance.
  • 19 November @ 2:59 PM - Maintenance completed. Services restored.  
  • 19 November @ 2:50astro email services will be taken off-line to adjust performance tuning parameters. email services are expected to be back on-line within 20 minutes.

Emergency Maintenance

  • 3 November @ 02:14 PM - DocuShare is back online in WRITE mode. Log files indicate that our backups are running without any problems.
  • 2 November @ 9:37 - The DocuShare server,, has been recovered via the indomitable and resolute efforts of Paul Morris and Patrick Goetz (CNS Systems), to whom we owe prodigious gratitude for their invaluable work. DocuShare is currently online in READ-ONLY mode while we configure, initiate and test a cron based backup to cerberus, our Oracle RAID.
  • 23 October @ 7:00 AM - was taken offline, it's internal disks removed and copies were made. Booting from a Redhat DVD, the two mirrored disks were reviewed and found to be inconsistent. Recover efforts continue. DocuShare will be offline for an indeterminate period of time.
  • 22 October @ 01:42 PM - DocuShare server database crashed. Emergency maintenance in progress.
  •  """""""""""@ 02:10 PM -, DocuShare server rebooted.
  • """"""""""" @03:18 PM - Server attempted to run a disk check and hung at 11.4%. Rebooted and it hung again at 11.4% during the disk check. Will allow the process to run overnight. And re-evaluate in the morning.


Scheduled Maintenance

  • 16 November @ 6:00 PM - Maintenance completed.  All services restored.
  • 16 November @ 5:00 PM - Services Server Upgrade.
  • 13 October @ 7:15 AM - Maintenance completed. All services restored.
  • 13 October @ 7:00 AM - gemini zones shutdown; gemini rebooted.
  • 13 October @ 6:30 AM - Security updates need to be applied to gemini, astro, ftp, and bigbubba. The Updates are already installed but the systems require rebooting for the updates to be applied. The following services will be off-line until the reboots are completed and services restarted:


Unscheduled Maintenance


  • astro email services will be taken off-line at 2:50 PM today to adjust performance tuning parameters. email services are expected to be back on-line within 20 minutes.
  • 17 September @ 9:49 PM Mail services restored. Maintenance completed. 
  • 17 September @ 8:15 PM Mail services will be disabled tor re-index the director account.
  • 22 August @ 12:54 PM Mail services restored. Maintenance completed. 
  • 22 August @12:45 PM Stop, Refresh and Restart mail services to clear out zombie processes. 

Planned Maintenance


  • 24 July @10:48am Resuming normal operations
  • 24 july @7:00am -12:00pm all major service will be offline for some period during this maintenance window
  • 15 July @ 12:45 PM - astro email  back online.
  • 15 July @ 12:15 PM - astro email (& thus webmail) SSL Security Certificates Update Click on Core Services Status above for grisly details
  • 13 July @ 5:00 PM - astro email (& thus webmail) SSL Security Certificates Update: :Click on Core Services Status above for grisly details.

    The SSL security certificates must be updated for mail services on astro to provide compatibility with mail services. SSL Certificate configuration continues to be a recondite undertaking with a plethora of convoluted methodologies and a poverty of guidance. In short, this may take awhile.

    5:00 PM astro email and webmail will be taken off-line for the SSL certificate updates.

  • 1 May 2015 @ 11:30  - In preparation for upcoming rewiring of our switch closet in RLM, there will be work performed to clear off a switch to use in the refit.  What this means for the user base is that there is a possibility that the connections in your rooms may be momentarily lost as connections are rerouted.  This is a joint project with CNS and is needed to get the refit that the closet desperately needs.  If you experience any lasting problems with your computer or phone connections, please send an email to helpdesk.  None of our core services are planned to be down due to this maintenance.  This only effects connections to individual offices.
  • 13 May @ 5:30 PM - astro and our other Solaris servers, ftp and gemini, will be rebooted at 5:30 PM today to apply maintenance and security patches. The updates have already been installed but require a reboot to be applied. The following services will be affected and be offline while the system are rebooting which hopefully will take less than 15 minutes.
    • astro email
    • webmail
    • Anonymous FTP
  • No labels